Four numbers have a median of 30, a mode of 50 and a mean of 29. What are the numbers?


1
Expert's answer
2021-01-14T17:24:09-0500

Four numbers: a, b, c, d.

A mode of 50, so at least two values will be 50.

Let a=50, b=50

A median of 30, so

"\frac{b+c}{2}=30 \\

b + c = 60 \\

50 + c = 60 \\

c = 10"

A mean of 29, so

"\frac{a+b+c+d}{4}=29 \\

a+b+c+d = 116 \\

50 + 50 + 10 + d = 116 \\

d = 6"

Answer: 50, 50, 10, 6.